OpenStudy (anonymous):

Find the slope of the line 5x-4y= -2

OpenStudy (dumbcow):

get y by itself first

OpenStudy (anonymous):

ok how do i go about doing so im stumnped with slope stuff

OpenStudy (anonymous):

step 1 -4y = -2 - 5x step 2 4y = 2 + 5x step 2 y = 5/4 x + 2/4 hence your slope is 5/4

OpenStudy (dumbcow):

thats how ;)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np